OpenStudy (anonymous):

Find the rectangular coordinates of the point with the polar coordinates. (3, 2pi/3)

OpenStudy (cggurumanjunath):

x=r cos theta ; y = r sin theta

OpenStudy (cggurumanjunath):

r = 3

OpenStudy (cggurumanjunath):

theta = (2pi/3)

OpenStudy (shamim):

x=? y=?

OpenStudy (cggurumanjunath):

x= 3 cos (2pi/3) ; y = 3 sin (2pi/3)

OpenStudy (cggurumanjunath):

(x,y) = 3 *cos 120 ; y = 3*

OpenStudy (cggurumanjunath):

y=3*sin 120

OpenStudy (shamim):

Cos120=? @katiexo

OpenStudy (shamim):

Sin120=?

OpenStudy (anonymous):

cos of 120 = .8142? sin 120 = .5806?

OpenStudy (anonymous):

i'm in radians

OpenStudy (shamim):

Cos120=cos(90+30)=?

OpenStudy (anonymous):

-1/2

OpenStudy (shamim):

x=?

OpenStudy (anonymous):

3

OpenStudy (shamim):

No

OpenStudy (anonymous):

-3/2

OpenStudy (anonymous):

wait x=-3/2 and since sin(120) is sqrt of 3/2 then y = 3*sqrt of 3/2? @shamim

OpenStudy (shamim):

Correct!!!!

OpenStudy (shamim):

Thanks!

OpenStudy (shamim):

Anyway i m in bus travelling to my college. So late to reply